# Little By Little: Continual Learning via Self-Activated Sparse Mixture-of-Rank Adaptive Learning

### v2

@@ A.9 Statistical Analyses of Contributing Rank Activations
1. Sparse Mixture: only a small subset of ranks is needed. Across all layers and positions, we find that fewer than 10% of the total ranks suffice to capture 99% of the activations. This highlights the extreme sparsity of MoRA’s self-activated mixture: most ranks remain dormant for any given input, while a compact set of highly relevant ranks drives the adaptation.
@@
To provide a broader view, Fig. 8 shows the required rank counts for every module in the pre-trained model. We observe that most attention modules requires around 6–12 ranks, while the second MLP projection generally demands more ranks in early layers, peaking in the first few blocks, and then steadily declines in deeper layers.
Coupling the rank activation budget with rank pruning, MoRA adapts the number of ranks needed to activate at each layer and module. This adaptive sparsity maximizes the efficient use of newly acquired knowledge during continual learning.

### v3

@@ A.13 Statistical Analyses of Contributing Rank Activations
1. Sparse Mixture: only a small subset of ranks is needed. Across all layers and positions, we find that fewer than 10% of the total ranks suffice to capture 99% of the activations. This highlights the extreme sparsity of MoRAM ’s self-activated mixture: most ranks remain dormant for any given input, while a compact set of highly relevant ranks drives the adaptation.
@@
To provide a broader view, Fig. 9 shows the required rank counts for every module in the pre-trained model. We observe that most attention modules require around 6–12 ranks, while the second MLP projection generally demands more ranks in early layers, peaking in the first few blocks, and then steadily declines in deeper layers.
Coupling the rank activation budget with rank pruning, MoRAM adapts the number of ranks needed to activate at each layer and module. This adaptive sparsity maximizes the efficient use of newly acquired knowledge during continual learning.